# Self-Rewarding Vision-Language Model via Reasoning Decomposition

### v1

@@ 2.3 Theoretical Analysis
We analyze why incorporating a perception-level reward improves vision-language RL training compared to relying solely on answer rewards. In standard RL training, the objective is driven only by the correctness of the final answer $a$:
$\nabla_{\theta}\,\mathbb{E}_{s\sim\pi_{\theta}}[r_{\text{ans}}(a,a^{*})],$ (5)
where $s=(t,a)$ is the trajectory with reasoning trace $t$ and final answer $a$. Since $r_{\text{ans}}$ depends solely on the match between $a$ and the ground truth $a^{*}$, the intermediate reasoning process $t$ receives no direct supervision. For VLMs, this is particularly problematic: the LLM backbone is far stronger than the vision encoder, and thus dominates the generation of $t$. With continued RL training under a single reward, gradients become high-variance and the model learns to exploit language priors – achieving correct answers without grounding in the visual input.
Reward decomposition.
We instead optimize a joint objective
$J(\theta)\;=\;\mathbb{E}_{s\sim\pi_{\theta}}\!\left[\,r_{\text{visual}}(c,x)\;+\;\,r_{\text{ans}}(a,a^{*})\right],$ (6)
where $r_{\text{visual}}$ measures informativeness of the visual perception caption $c$ of visual input $x$, and $r_{\text{ans}}$ is the final answer reward. The resulting gradient
$\nabla_{\theta}J(\theta)\;=\;\alpha\,\nabla_{\theta}r_{\text{cap}}\;+\;\beta\,\nabla_{\theta}r_{\text{ans}}$ (7)
anchors updates to both visual perception and language reasoning modules, providing direct signals for both perception and language reasoning.
From an information-theoretic perspective, let $I$ denote the visual input, $Q$ the question, $C$ the visual perception representation, and $A$ the final answer. Mutual information $I(U;V)$ measures how much knowing $U$ reduces uncertainty about $V$ (Shannon, 1948). If training relies only on $r_{\text{ans}}$, the model primarily maximizes I(A;Q)I(A;Q)—making answers strongly dependent on the question—while neglecting $I(A;I)$, the dependence of answers on the visual input. This permits shortcut solutions that bypass perception. By additionally optimizing $r_{\text{cap}}$, which enforces high $I(C;I)$, the model strengthens the path from $I$ to $A$, thereby increasing $I(A;I)$ and ensuring that answers remain grounded in visual perception rather than language-only correlations.
@@ 2.4 Data Preparation
Figure 2: We prompt Qwen-2.5-VL-7B to create the SFT cold-start dataset to learn the ideal format to get a quick start in the RL stage to reduce training steps and training time. Our cold start SFT dataset is the intersection of three source datasets, creating a final set of 9K examples. Our filtration process in Sec 2.4 guarantees zero false positives across text-only, caption-based, and direct CoT reasoning.
[…]
@@ 3 Experiments
To implement our Vision-SR1, we employ Qwen-2.5-VL-3B and 7B as base models. We use SFT to train our the base model on the SFT cold start dataset to learn the ideal format. Following the initial SFT stage, we train the model with GRPO. The RL phase is trained for 1 epoch on the Vision-SR1-47K dataset. During training, the policy model first generates visual perceptions from the input image, then produces language reasoning and final answer. Before computing the final GRPO advantage, we compute a self-reward for visual perception. This reward is derived from the policy model’s ability to answer the question using only its own generated visual perceptions, without relying on the input image $i$.
@@ 3.1 Baseline Methods
Vision-R1 (Huang et al., 2025): The first R1-style reinforcement learning approach, which relies solely on answer rewards as the training signal. However, since the original Vision-R1 was trained only on math-domain data and performs poorly on general-domain reasoning, we reproduce it using our 47K dataset to ensure a fair comparison.
Perception-R1 (Xiao et al., 2025): Similar in training style to Vision-R1, but incorporates pre-extracted visual annotations as an additional reward signal. These visual annotations are derived from a state-of-the-art proprietary multimodal LLM (not specified in the paper).
@@ 3.2 Benchmarks and Metrics
General Visual Understanding. We evaluate general visual understanding across five diverse benchmarks. MMMU (Yue et al., 2024) tests cross-modal reasoning and subject knowledge with 11.5K college-level, four-choice questions spanning six disciplines. MMMU-Pro (Yue et al., 2025) increases the difficulty with ten choices per question and adds a challenging vision-only setting, where all text is embedded within the image to necessitate robust visual parsing. MM-Vet (Yu et al., 2024b) assesses a broad range of integrated vision-language skills – including recognition, OCR, and math – using a unified LLM-based evaluation score. RealWorldQA (xAI, 2024) features $\sim$700 real-world images from vehicle captures, paired with spatially grounded questions that require verifiable answers. VisNumBench (Weng et al., 2025) specifically targets visual number sense through $\sim$1.9K questions covering seven numerical attributes and four estimation tasks.
@@
Hallucination Diagnosis. To diagnose model failures, we use HallusionBench (Guan et al., 2024), a benchmark designed to pinpoint specific errors: (i) language-side hallucination, where visual context is ignored, and (ii) visual-illusion errors, where the image is misinterpreted. The benchmark’s binary yes/no format enables precise error analysis.
For our evaluations, we all use Gemini-2.5-flash (Comanici et al., 2025) to judge response correctness on non-multiple choice format question, serving as a proxy for human judgment.
@@ 3.3 Experimental Results
Methods MMMU MMMU $M$ RealWorld VisNum Math MATH Hallusion Avg.
-Pro -Vet QA Bench Verse -Vision Bench
Visionary-R1 (3B) by Xia et al. (2025) 27.4 30.6 63.8 56.9 10.0 45.0 36.5 30.0 37.5
Percention-R1 (7B) by Xiao et al. (2025) 36.8 40.9 78.4 69.4 15.9 57.6 41.2 65.4 50.7
Vision-R1 (7B) by Huang et al. (2025) 34.9 42.8 70.2 60.1 33.0 57.3 42.9 65.2 50.8
@@
Zero-shot Inference (before RL) 30.5 25.5 68.4 65.4 15.7 44.3 35.8 40.9 40.8
Supervised Fine-tuning (before RL) 39.5 49.4 70.2 62.1 29.3 47.8 39.9 68.1 50.8
Vision-R1 47K data (fair comparison) 40.3 49.5 63.3 63.0 36.7 45.5 38.8 67.4 50.6
Vision-SR1 (our method) 40.8 49.6 69.7 66.1 41.9 48.5 38.5 68.3 52.9
@@
Zero-shot Inference (before RL) 34.2 33.5 67.1 68.5 21.4 49.2 41.1 51.7 45.8
Supervised Fine-tuning (before RL) 41.8 51.8 79.4 65.6 35.7 55.5 42.8 67.8 55.1
Vision-R1 47K data (fair comparison) 47.7 54.8 78.9 69.9 39.4 54.7 46.0 67.6 57.4
Vision-SR1 (our method) 49.1 57.2 76.2 71.6 42.6 56.5 46.7 69.8 58.8
@@
Table 2 presents a comprehensive comparison of Vision-SR1 with several baseline methods across diverse vision-language benchmarks. For example, with the Qwen2.5-VL-7B backbone, Vision-SR1 reaches 49.1 on MMMU-Pro and 57.2 on MMMU, outperforming Vision-R1 fair comparison runs (47.7 and 54.8, respectively). When averaged across all benchmarks, Vision-SR1 establishes a clear margin of improvement. With the 7B backbone, it achieves an average score of 58.8, compared to 57.4 for Vision-R1 and 55.1 for supervised fine-tuning. Even with the smaller 3B backbone, Vision-SR1 attains 52.9 average, outperforming all comparable baselines. These results demonstrate that Vision-SR1 significantly outperforms prior baseline models across both general-purpose and math-specific visual reasoning tasks, validating the effectiveness of our approach.
3.3.2 Ablation Study on Self-Reward
We train a control version of our model without the visual perception self-reward (Vision-SR1 w/o self-reward). This ablated model still follows a structured output (visual perception, CoT reasoning, and answer) but is optimized only with answer reward and format rewards. The self-visual reward for self-evaluating visual perception is removed. We note the only difference between Vision-SR1 w/o self-reward and Vision-R1 (Huang et al., 2025) lies in the output structure, i.e., using different system prompts, while all supervision signals (answer reward and format rewards) remain the same. Interestingly, our system prompt yields slightly better performance (+0.3 on average). Table 3 reports the ablation results. We find that removing the visual perception reward consistently leads to worse overall performance compared to including self-reward in the training process.
Table 3: Results of ablation study: Vision-SR1 v.s. Vision-SR1 without visual perception self-reward.
@@ 3.3.2 Ablation Study on Self-Reward
Methods MMMU MMMU $M$ RealWorld VisNum Math MATH Hallusion Avg.
-Pro -Vet QA Bench Verse -Vision Bench
Vision-SR1 (3B) 40.8 49.6 69.7 66.1 41.9 48.5 38.5 68.3 52.9
  $\vdash$ w/o self-reward 40.0 48.0 67.4 62.6 41.6 47.7 38.9 65.8 51.5
Vision-SR1 (7B) 49.1 57.2 76.2 71.6 42.6 56.5 46.7 69.8 58.8
  $\vdash$ w/o self-reward 48.8 55.3 78.4 70.9 41.4 54.8 45.3 66.4 57.7
3.3.3 Analysis on Language Shortcut
Table 4: Language Shortcut Rate (LSR) across different benchmarks. Lower values indicate better performance, as a reduced LSR reflects fewer language shortcuts during reasoning.
General Visual Understanding Visual Math & Hallucination
Methods MMMU MMMU $M$ RealWorld VisNum Math MATH Hallusion Avg.
-Pro -Vet QA Bench Verse -Vision Bench
Vision-SR1 (3B) 6.49 9.0 6.4 0.07 5.4 10.3 8.3 0.10 5.6
  $\vdash$ w/o self-reward 7.36 8.0 5.1 0.07 4.2 11.4 9.2 0.09 7.6
Vision-SR1 (7B) 9.12 8.0 8.3 0.07 10.9 4.7 13.2 0.07 6.7
  $\vdash$ w/o self-reward 8.86 7.0 7.3 0.08 10.2 15.5 13.4 0.09 7.9
@@ 3.3.3 Analysis on Language Shortcut
Our evaluation, follows a two-step process and uses Gemini-2.5-flash as a judge: (1) Visual Perception Extraction: for each model output, we extracted the generated visual perception, denoted as $\hat{C}$. (2) Self-Containment Check: we then provide the $\hat{C}$ and the original question $Q$ to Gemini-2.5-Flash evaluator. If the evaluator can reproduce the correct ground-truth answer using only this information, $\hat{C}$ is deemed self-contained. Based on this process, we define the metrics: The Language Shortcut Rate (LSR) is defined as the percentage of instances where the model produces an incorrect (not self-contained) visual perception but still gives the correct final answer:
@@
$\text{LSR}=\frac{\#\{\text{incorrect visual perception{} \& correct answer}\}}{\#\{\text{total samples}\}}$
@@
A higher LSR indicates that the model is answering correctly while bypassing visual perception, suggesting reliance on language prior shortcuts. An LSR of 0 indicates no shortcutting, i.e., every correct answer is supported by a correct, self-contained visual perception.
@@
We compute the LSR for 7B model w/ and w/o self rewards on seven selected benchmarks for demo example in Table 4. An important finding is that the visual shortcut is the highest in multimodal mathematical reasoning, which raises important question to previous work R1-VL (Zhang et al., 2025), VLM-R1 (Shen et al., 2025), Vision-R1 (Huang et al., 2025): is multimodal RL training on mathematical datasets truly improving VLMs’ abilities to perform visual reasoning, or simply awakes the models’ language reasoning ability to guess without actually looking at visual information?
3.3.4 Analysis on Text-only Reasoning

### v2

@@ 2.4 Theoretical Analysis
We analyze why Multi-Reward Policy Optimization with separate advantage computation could improve VLM RL training compared to using only answer rewards. In standard RL training, the objective depends solely on final answer correctness:
$\nabla_{\theta}\,\mathbb{E}_{s\sim\pi_{\theta}}[r_{\text{ans}}(a,a^{*})]$ (10)
where $s=(t,a)$ contains visual reasoning and language reasoning trace $t$ and final answer $a$. Since $r_{\text{ans}}$ only measures whether $a$ matches ground truth $a^{*}$, the intermediate visual reasoning $t$ receives no direct supervision signal. For VLMs, the stronger LLM backbone dominates generation of $t$, and continued RL training leads to potential reward hacking where the model exploits language priors to achieve correct answers without visual grounding Pantazopoulos and Özyiğit (2025).
Multi-Reward Loss Decomposition. We decompose the loss computation itself into separate components as shown in Equation 9, where the actor loss handles visual and answer components separately (with the $KL$ regularization term following similar component-wise structure):
$\mathcal{L}_{\text{actor}}=-\lambda_{\text{ans}}\mathbb{E}[A_{\text{ans}}\log\pi_{\theta}(a_{\text{ans}})]-\lambda_{\text{visual}}\mathbb{E}[A_{\text{visual}}\log\pi_{\theta}(a_{\text{visual}})]$ (11)
Since each advantage is computed from different reward components (visual and answer, in Equation 6), this approach creates clear gradient paths from each reward to its corresponding components, enabling independent optimization of visual reasoning and language reasoning capabilities (Zhu et al., 2025; Lyu et al., 2025; Liu et al., 2026).
2.4.1 Computational Efficiency Analysis
A natural question with two-stage rollout training is whether it doubles the computational cost relative to standard one-stage GRPO. This section’s analysis shows two-stage rollout is only 10-15% more expensive than standard GRPO, while requiring no extra GPU computation.
In training, both rollout stages reuse the same VLM without loading additional models nor calling external APIs. In practice, two-stage rollout training adds only a modest overhead over standard GRPO. For example, training for 20 steps (per-device batch size 8, 8 GPUs) on a 7B model takes standard GRPO approximately 10.5 hours, while our two-stage training takes approximately 13 hours, an overhead of roughly 20%. Additionally, we compare against two alternative external visual reward strategies. 1): Using proprietary models (GPT-5 Singh et al. (2025), Claude Anthropic , Gemini Peng et al. (2025b)) as an external judge requires evaluating $N\times B$ responses per step, where $N$ is the number of rollouts per sample and $B$ is the effective batch size, which can cause API rate limits are hit quickly, pushing training time more than doubling the cost. Using a local open-source judge instead requires dedicating at least one GPU to inference, reducing training parallelism from $N$ to N1N-1 GPUs and introducing additional inference latency. By contrast, self-reward training avoids both penalties, making two-stage training practical and scalable without incurring the overhead of external judges.
@@ 3 Experiments
To implement our Vision-SR1, we use Qwen2.5-VL-3B and 7B, Mimo-7B-VL Team et al. (2025) as base models. We train the base model with GRPO. The RL phase is trained for 200 steps on the Vision-SR1-47K dataset. During training, the policy model first generates visual reasoning from the input image, then produces language reasoning and final answer. We then compute a self-reward for visual reasoning by re-prompting the frozen policy model to answer the question using only its generated visual reasoning, without access to the original image $i$. Finally, we compute advantages and log probabilities separately for each reward component and combine them in the final loss (Figure 1). [^2] The policy model remains frozen during both rollouts.
@@ 3.1 Baseline Methods
Vision-R1 (Huang et al., 2025b): The first R1-style reinforcement learning approach, which relies solely on answer rewards as the training signal. However, since the original Vision-R1 was trained only on the math domain and falters on general-domain reasoning, we reproduce it using our 47K dataset to ensure a fair comparison.
Perception-R1 (Xiao et al., 2025)’s training resembles Vision-R1 but incorporates pre-extracted visual annotations as an additional reward signal. These visual annotations are derived from a state-of-the-art proprietary multimodal LLM (not specified in the paper).
@@ 3.2 Benchmarks and Metrics
General Visual Understanding. We evaluate general visual understanding across five diverse benchmarks. MMMU (Yue et al., 2024) tests cross-modal reasoning and subject knowledge with 11.5K college-level, four-choice questions spanning six disciplines. MMMU-Pro (Yue et al., 2025) increases the difficulty with ten choices per question and adds a challenging vision-only setting, where all text is embedded within the image to necessitate robust visual parsing. RealWorldQA (xAI, 2024) features $\sim$700 real-world images from vehicle captures, paired with spatially grounded questions that require verifiable answers. VisNumBench (Weng et al., 2025) specifically targets visual number sense through $\sim$1.9K questions covering seven numerical attributes and four estimation tasks.
@@
Hallucination Diagnosis. To diagnose model failures, we use HallusionBench (Guan et al., 2024), a benchmark designed to pinpoint specific errors: (i) language-side hallucination, where visual context is ignored, and (ii) visual-illusion errors, where the image is misinterpreted. Because the benchmark uses binary yes/no questions with unambiguous ground truth, each incorrect response can be cleanly attributed to one of these two failure modes.
For our evaluations, we all use Gemini-2.5-flash (Comanici et al., 2025) to judge response correctness on non-multiple choice format questions, serving as a proxy for human judgment.
@@ 3.3 Experimental Results
Methods MMMU MMMU RealWorld VisNum Math MATH Hallusion Avg.
-Pro QA Bench Verse -Vision Bench
Visionary-R1 (3B) by Xia et al. (2025) 27.4 30.6 56.9 10.0 45.0 40.4 26.7 33.9
Percention-R1 (7B) by Xiao et al. (2025) 36.8 40.9 69.4 15.9 52.1 35.7 65.4 45.2
Vision-R1 (7B) by Huang et al. (2025b) 34.9 42.8 60.1 33.0 57.3 51.2 32.2 44.5
@@
Zero-shot Inference (before RL) 30.5 25.5 65.4 15.7 44.3 40.4 27.1 35.5
Vision-R1 47K data (fair comparison) 40.3 49.5 63.0 36.7 42.8 29.9 67.4 47.1
Vision-SR1 (ours) 40.8 49.6 66.1 41.9 45.8 29.3 68.3 48.8
@@
Zero-shot Inference (before RL) 34.2 33.5 68.5 21.4 49.2 31.9 51.7 41.5
Vision-R1 47K data (fair comparison) 39.8 51.8 66.6 43 53.2 33.8 66.6 50.7
Vision-SR1 (ours) 40.7 52.2 69.2 43.5 54.5 36.2 68.9 52.2
Backbone model: Mimo-VL-7B
Zero-shot Inference (before RL) 38.0 45.6 68.2 30.2 35.5 21.6 71.9 44.4
Vision-R1 47K data (fair comparison) 38.7 47.3 67.1 33.5 35.3 25.7 74.3 46.0
Vision-SR1 (ours) 39.3 49.5 68.1 44.6 40.0 29.6 75.6 49.5
@@
Table 2 compares Vision-SR1 with several baseline methods across diverse vision-language benchmarks. With the Qwen2.5VL-7B backbone, Vision-SR1 reaches 40.7 on MMMU-Pro and 52.2 on MMMU, outperforming Vision-R1 fair comparison runs (34.9 and 42.8, respectively). When averaged across all benchmarks, Vision-SR1 establishes a clear margin of improvement. With the 72B backbone, it achieves an average score of 52.2, compared to 44.5 for Vision-R1. Even with the smaller 3B backbone, Vision-SR1 achieves 48.8 average, outperforming all comparable baselines. For results on Mimi-VL-7B, a model outside the Qwen-VL family, we observe a similar trend: the average accuracy improves from 44.4 to 49.5. This shows that our method generalizes beyond the Qwen-VL. These results demonstrate that Vision-SR1 outperforms prior baseline models across both general-purpose and math-specific visual reasoning tasks, validating the effectiveness of our approach.
Table 3: Our method also can improve VLMs’ abilities on spatial reasoning and language shortcut (LS) robustness.
Methods ViLP MMSI Omni Avg.
(LS) -Bench Spatial
Backbone: Mimo-VL-7B
before RL 56.4 28.2 40.3 41.6
Vision-R1 58.2 27.7 40.4 42.1
Vision-SR1 59.3 28.0 42.7 43.3
Backbone: Qwen2.5-VL-7B
before RL 45.1 24.0 27.3 32.1
Vision-R1 51.3 21.9 31.1 34.8
Vision-SR1 52.6 27.7 44.2 41.5
3.3.2 Ablation study on Spatial Reasoning and Language Shortcut Datasets
In addition to evaluating Vision-SR1 on standard visual-reasoning benchmarks, we further evaluate its effectiveness on additional datasets to probe two complementary challenges: spatial reasoning and language-shortcut (LS) robustness. MMSI-Bench Yang et al. (2025) and OmniSpatial Jia et al. (2025) target multi-image spatial understanding, requiring models to integrate spatial relationships across multiple images. In contrast, ViLP Luo et al. (2025) evaluates language shortcuts by pairing each question with images that can be answered either through textual priors alone or only through pure visual reasoning. Table 3 shows that Vision-SR1 generalizes well to spatial reasoning benchmarks and substantially improves robustness to visual–language shortcuts. In particular, explicitly generating visual descriptions helps the model avoid shortcut behavior and rely more on the actual visual content. Next we propose a systematic way to evaluate VLMs’ language shortcut frequencies on standard VLM benchmarks.
3.3.3 Analysis on Language Shortcut
Table 4: Language Shortcut Rate (LSR) across different benchmarks. Lower values indicate better performance, as a reduced LSR reflects fewer language shortcuts during reasoning. Adding additional reward supervision can reduce the change of visual reasoning reward hacking.
@@ 3.3.3 Analysis on Language Shortcut
Methods MMMU MMMU RealWorld VisNum Math MATH Hallusion Avg.
-Pro QA Bench Verse -Vision Bench
Vision-SR1 (3B) 7.5 6.3 10.8 5.4 10.3 8.3 10.1 9.4
  $\vdash$ w/o self-reward 9.0 9.6 11.9 4.2 11.4 9.2 8.5 10.4
Vision-SR1 (7B) 8.0 6.5 13.4 4.2 11.5 10.7 6.8 9.8
  $\vdash$ w/o self-reward 8.7 5.3 10.8 3.9 12.7 10.7 9.1 10.1
@@
Our two-step evaluation uses Gemini-2.5-flash as a judge: (1) Visual Perception Extraction: for each model output, we extracted the generated visual reasoning, denoted as $\hat{C}$. (2) Self-Containment Check: we then provide the $\hat{C}$ and the original question $Q$ to Gemini-2.5-Flash evaluator. If the evaluator can reproduce the correct ground-truth answer using only this information, $\hat{C}$ is self-contained. Based on this process, we define the Language Shortcut Rate (LSR) as the percentage of instances where the model produces an incorrect (not self-contained) visual reasoning but still gives the correct final answer:
@@
$\text{LSR}=\frac{\#\{\text{incorrect visual reasoning{} \& correct answer}\}}{\#\{\text{total samples}\}}$
@@
A higher LSR indicates that the model is answering correctly while bypassing visual perception, suggesting reliance on language prior shortcuts. An LSR of 0 indicates no shortcutting, i.e., every correct answer is supported by a correct, self-contained visual reasoning.
@@
We compute the LSR for 7B model w/ and w/o self rewards on seven selected benchmarks for demo example in Table 4. Visual shortcuts pervade multimodal mathematical reasoning, which raises important questions in previous work R1-VL (Zhang et al., 2025), VLM-R1 (Shen et al., 2025), Vision-R1 (Huang et al., 2025b): is multimodal RL training truly improving VLMs’ abilities to perform visual reasoning or simply awakening the models’ language reasoning ability to guess without actually looking at visual information?
3.3.4 Analysis of Visual Attention Change
Figure 2: Post-training encourages the ViT to devote more attention to visual tokens at both early feature extraction and late integration stages, while compressing intermediate processing.
We analyze layer-wise shifts in visual attention between the base model and Vision-SR1 to understand how post-training affects visual processing. We sample 1,000 images from the ViLP dataset and compute, for each image, the per-layer difference in visual attention weights between the post-trained and pretrained models via L2 distance. Averaging these differences across all 50 samples yields a layer-wise breakdown of attention redistribution over visual tokens. As shown in Figure 2, the changes cluster into two distinct patterns. In the early layers (0–7), post-training increases visual attention, with the largest gain at Layer 6 (+10.2%), showing the model learns to extract richer low-level visual features earlier in the network. The middle layers’ (8–13) decrease indicates a redistribution rather than a uniform increase in attention. In the late layers (14–27), visual attention rises again, peaking at Layer 20 (+9.2%), which shows enhanced visual re-engagement during the visual integration and output generation stages.
4 Related Work
4.1 Post-Training Vision-Language Models
Recent vision-language models have increasingly leveraged post-training alignment techniques, including instruction tuning and reinforcement learning, to enhance general-purpose multimodal performance (Liu et al., 2023b; Bai et al., 2025; Chen et al., 2024; et al, 2024; Huang et al., 2025b). For example, LLaVA (Liu et al., 2023b) is tuned on GPT-4 generated (image, question, answer) pairs, coupling a CLIP encoder with Vicuna to produce a visual chat assistant that imitates some GPT-4 vision capabilities. InstructBLIP (Dai et al., 2023) introduces an instruction-aware query transformer tuned on 26 datasets, which yields a model that substantially outperforms even larger models on zero-shot benchmarks. Beyond standard instruction-tuning methods like LLaVA and InstructBLIP, recent work increasingly uses reinforcement learning (RL) to align vision-language models for better reasoning (Huang et al., 2025b; Xia et al., 2025; Xiao et al., 2025). Many of these methods, inspired by techniques from DeepSeek-R1 (DeepSeek-AI et al., 2025), focus on sophisticated reward engineering. Strategies include providing step-wise rewards to supervise the intermediate reasoning (Zhang et al., 2025), adding explicit visual annotations to ground truth for calculating visual rewards (Xiao et al., 2025), and applying RL in a two-stage curriculum that first strengthens text-only reasoning (Peng et al., 2025b). As a complementary approach, RL from AI Feedback for VLMs demonstrates that preference-based alignment is also a powerful signal, showing it can substantially reduce object hallucination by learning from AI-generated feedback (Yu et al., 2024).
4.2 Self-Rewarding Reinforcement Learning
The existing reinforcement learning with verifiable rewards (RLVR) methods heavily rely on high-quality reward models or human feedback, creating a major bottleneck for scalability (Peng et al., 2025a; Dai et al., 2025; Li et al., 2025c; Luu et al., 2025). To overcome this, recent work explores self-rewarding approaches, where the model itself provides intrinsic reward signals during RL post-training, an idea first pioneered by Yuan et al. (2025). Building on self-rewarding language models, methods replace external reward models with the model’s own confidence and uncertainty (logit-based self-certainty) or self-verification of its solutions, and even elicit a latent endogenous reward already present inside base LLMs (Zhao et al., 2025; Li et al., 2025a; Simonds et al., 2025; Zheng et al., 2025; van Niekerk et al., 2025; Huang et al., 2025a; Zhou et al., 2025). For example, RLIF leverages self-certainty as a reward, achieving comparable performance to GRPO while improving out-of-distribution generalization (Zhao et al., 2025). Similarly, RLSC optimizes a self-confidence reward to secure large accuracy gains with only a few training samples (Li et al., 2025a).
Although self-generated reward signals have thrived in text-only LLMs, only a few works extend this idea to VLMs (Zhou et al., 2024; Lee et al., 2025; Holmes and Chi, 2025), largely due to the complexity of the visual modality and the difficulty of properly defining and evaluating reward signals that capture visual perception. Recent progress includes Calibrated Self-Rewarding, which iteratively generates candidates, self-scores them with step-wise, visually constrained rewards, and fine-tunes via direct preference optimization (DPO) (Zhou et al., 2024). Similarly, RG-VLM uses a VLM to directly label rewards for offline trajectories in long-horizon visual tasks, serving as an auxiliary signal that boosts generalization (Lee et al., 2025). Beyond judgment-based signals, ARES derives dense shaped rewards from attention weights to accelerate learning under sparse or delayed feedback (Holmes and Chi, 2025). These works show that internal visual signals can provide rich reward feedback for VLM alignment without costly supervision, yet the reward is not integrated end-to-end, where the policy receives both visual perception and answer rewards during training.
5 Conclusion and Future Work
We introduce Vision-SR1, a self-rewarded reinforcement learning framework that decomposes vision-language understanding into visual reasoning and language reasoning components. Our approach uses the VLM itself to generate explicit rewards for visual understanding, then applies Multi-Reward Policy Optimization to provide clear gradient attribution and backpropagation pathways for each reward component. Vision-SR1 strengthens visual perception and reduces language shortcuts, thereby improving VLM performance across several domains of vision-language tasks. Our proposed metric LSR further shows how perception reward lowers the tendency of models to answer via language shortcut rather than genuine visual reasoning.
This work opens up several future research directions. First, future work can focus on improving the efficiency of the visual reasoning then think generation format by treating the visual reasoning component as latent thinking, thereby reducing the number of decoded tokens while still enabling reward attribution to latent visual processes during the RL phase. It is also important to recognize that some of the observed mathematical gains from RL training in VLMs may come from spurious effects – for instance, recalibrating the LLM backbone’s output distribution can boost multimodal math performance without true visual grounding (Shao et al., 2025). This suggests that improvements in accuracy may sometimes reflect better exploitation of language shortcuts rather than genuine perception gains. Therefore, future work can also explore more analysis to disentangle visual grounding from shortcut learning.
Acknowledgments
Boyd-Graber is supported by nsf grant 2229885, the University of Maryland (umd) TRAILS (Trustworthy ai in Law & Society) initiative. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the researchers and do not necessarily reflect the views of the National Science Foundation.
Reproducibility Statement
To ensure the reproducibility of our research, we provide information regarding our prompt templates and experimental setup in the main paper and Appendix. All datasets and code will be released upon conference decision release.
Appendix A Appendix
A.1 The Use of Large Language Models (LLMs)
We acknowledge the use of large language models (LLMs) as assistive tools in this research. Our use of LLMs was limited to refine grammar and improve language clarity. All outputs from these models were meticulously reviewed, revised, and verified by the authors, who retain full responsibility for all content presented in this paper.
Appendix B Experiment Details
B.1 Prompt Templates
[…]
B.2 LLM-as-a-Judge Prompt
We use Gemini-2.5-flash as our LLM-as-a-Judge to evaluate
Question: {Question}
B.2.1 Analysis on Text-only Reasoning